Stealth Fitness Center issues 9%, 10-year bonds with a face amount of $100,000. The market interest rate for bonds of similar risk and maturity is 8%. Interest is paid semiannually. At what price will the bonds be issued?
Answer #1

When the coupon rate is greater than the market rate that means the bonds are issued at premium.

Issue price of bonds= Present value of principal+Present value of interest

= $45639+61156.485= $106795.485

Present value of principal= $100000*0.45639= $45639

Semiannual cash interest payment= $100000*9%*6/12= $4500

Present value of interest= $4500*13.59033= $61156.485

0.45639, is the Present value of $1, n= 20 years (10*2) and i= 4% (8%/2)

13.59033, is the Present value of ordinary annuity, n= 20 years (10*2) and i= 4% (8%/2)